# Robust Transmission Design for RIS-Assisted Integrated Sensing and Communication Systems

### v1

@@ Abstract
As a critical technology for next-generation communication networks, integrated sensing and communication (ISAC) aims to achieve the harmonious coexistence of communication and sensing. The degrees-of-freedom (DoF) of ISAC is limited due to multiple performance metrics used for communication and sensing. Reconfigurable Intelligent Surfaces (RIS) composed of metamaterials can enhance the DoF in the spatial domain of ISAC systems. However, the availability of perfect Channel State Information (CSI) is a prerequisite for the gain brought by RIS, which is not realistic in practical environments. Therefore, under the imperfect CSI condition, we propose a decomposition-based large deviation inequality approach to eliminate the impact of CSI error on communication rate and sensing Cramér-Rao bound (CRB). Then, an alternating optimization (AO) algorithm based on semi-definite relaxation (SDR) and gradient extrapolated majorization-maximization (GEMM) is proposed to solve the transmit beamforming and discrete RIS beamforming problems. We also analyze the complexity and convergence of the proposed algorithm. Simulation results show that the proposed algorithms can effectively eliminate the influence of CSI error and have good convergence performance. Notably, when CSI error exists, the gain brought by RIS will decrease with the increase of the number of RIS elements. Finally, we summarize and outline future research directions.
@@ I Introduction
Numerous emerging applications for future networks, including vehicle-to-everything, smart homes, and industrial Internet of thing (IoT), require high-quality communication performance and demand accurate sensing capabilities [1, 2]. The communication and sensing functionality can be simultaneously realized in a unified platform due to their similar hardware architecture and signal processing algorithms. Motivated by these considerations, research on integrated sensing and communication (ISAC) is currently underway to empower next-generation communication systems with sensing capabilities, such as mobile edge computing (MEC) [3], federated learning [4], and beamforming [5]. ISAC can realize dynamic spectrum sharing between sensing and communication, achieving harmonious co-existence between radar and communication systems [6].
Reconfigurable intelligent surface (RIS) is a planar surface consisting of a large number of low-cost and nearly passive reconfigurable elements that can be dynamically controlled to induce appropriate amplitude and phase shifts to incident signals [7, 8, 9, 10]. RIS has been identified as a promising technology to enhance the energy and spectral efficiency of communication systems and radar systems. Research on deploying RISs into communication systems mainly focuses on joint beamforming design [11, 12, 13] and robust transmission designs [14, 15]. On the other hand, research on deploying RISs into radar systems focuses on enhancing sensing coverage in environments characterized by rich scattering, such as joint beamforming works [16, 17, 18] and sensing architecture [19]. Additionally, RISs can be deployed in scenarios where infrastructures, such as the Global Positioning System (GPS), are insufficient to estimate the locations of UEs and scatterers. To be specific, theoretical bounds for localization have been investigated in [20], signal strength-based localization schemes have been proposed in [21], and time-of-arrival (TOA) and angle of departure (AOD) based methods have been investigated in [22, 23].
@@ I-A The Contribution of This Work
Our study focuses on robust beamforming for RIS-assisted ISAC systems. We propose a novel approximation method to approximate the probability constraints for the communication rate and the lower bound of the estimation accuracy, i.e., the Cramér-Rao bound (CRB). We then propose an alternating optimization (AO) algorithm to solve the robust beamforming problem, with two schemes for handling the non-convex discrete RIS phase shifts constraints. Our proposed robust beamforming method is demonstrated to approach the probability constraints tightly and effectively handle the discrete RIS phase shifts.
@@ II System Model
Figure 1: An RIS-assisted ISAC system.
The considered system is illustrated in Fig. 1, where a dual-functional radar and communication (DFRC) BS with $N$ antennas communicates with $K$ single-antenna user while sensing $L$ targets, with the aid of an $M$-element RIS. The transmit antennas are uniform linear arrays (ULAs). We assume that the inter-distance of the transmit antennas and RIS elements are half-wavelength. The set of RIS elements, communication users, and targets are denoted by $\mathcal{M}$, $\mathcal{K}$, and $\mathcal{L}$. The phase shifts of the RIS elements are controlled by an RIS controller, which communicates wirelessly with the DFRC BS for cooperative transmission. The phase matrix of the RIS is represented by $\boldsymbol{\Theta}\triangleq\text{diag}\left(\theta_{1},\cdots,\theta_{M}\right)$. We assume an ideal RIS reflection magnitude model where each element is fixed at $|\theta_{m}|=1,\forall m\in\mathcal{M}$, as well as a practical RIS reflection phase shifts model where each $\theta_{m}$ can only take $d$ finite values, equally spaced in $[0,2\pi)$. The discrete set of phase shifts for each RIS element is given by
@@ II-A Communication Signal Model
where $\boldsymbol{h}_{\text{BU},k}\in\mathbb{C}^{N\times 1}$ is the channel vector between the BS and the $k$-th user, $\boldsymbol{h}_{\text{RU},k}\in\mathbb{C}^{M\times 1}$ is the channel vector between the RIS and the $k$-th user, and $\boldsymbol{H}_{\text{BR}}\in\mathbb{C}^{N\times M}$ is the channel matrix between the BS and the RIS. The transmit beamforming matrix of the DFRC BS is $\boldsymbol{S}=\left[\boldsymbol{s}_{1},\boldsymbol{s}_{2},\cdots,\boldsymbol{s}_{K}\right]\in\mathbb{C}^{N\times K}$. The transmit signal matrix is $\boldsymbol{X}=\left[\boldsymbol{x}_{1},\boldsymbol{x}_{2},\cdots,\boldsymbol{x}_{K}\right]^{T}\in\mathbb{C}^{K\times T}$, where $T$ is the number of signal samples. The signal vectors in $\boldsymbol{X}$ are assumed to be orthogonal to each other, i.e., $\boldsymbol{X}\boldsymbol{X}^{H}=\textbf{I}_{K}$. Moreover, $\boldsymbol{z}_{k}$ is the additive white Gaussian noise (AWGN) at the $k$-th user, which is assumed to follow the circularly symmetric complex Gaussian (CSCG) distribution, i.e., $\boldsymbol{z}_{k}\sim\mathcal{CN}\left(0,\sigma_{\text{com},k}^{2}\textbf{I}\right)$ with $\sigma_{\text{com},k}^{2}$ denoting the variance of the AWGN.
Due to the passive characteristic of the RIS, the cascaded BS-RIS-user channel at the transmitter (CBRUT) is difficult to obtain. The direct channel state information at the transmitter (DCSIT) is also difficult to obtain in the complex wireless propagation environments. Hence, we assume the imperfect DCSIT and the imperfect CBRUT. The imperfect DBIUT can be expressed as
@@
where $\boldsymbol{H}_{\text{BRU},k}\triangleq\text{diag}\left(\boldsymbol{h}_{\text{RU},k}\right)^{H}\boldsymbol{H}_{\text{BR}}^{H}\in\mathbb{C}^{M\times N}$ is the cascaded BS-RIS-user channel. The CSI error, i.e., $\Delta\boldsymbol{h}_{\text{BU},k}$ and $\Delta\boldsymbol{H}_{\text{BRU},k}$ are all assumed to follow the CSCG distribution. The CSI error can be expressed as
@@ II-B Sensing Signal Model
The received sensing signal at DFRC BS is $\boldsymbol{Y}_{\text{sen}}$, which can be expressed as
@@
where $\boldsymbol{G}_{\text{sen},l}\triangleq\alpha_{l}\boldsymbol{\mathrm{a}}\left(\phi_{l}\right)\boldsymbol{\mathrm{a}}\left(\phi_{l}\right)^{T}\in\mathbb{C}^{N\times N}$ denotes the sensing response matrix, $\alpha_{l}$ is the reflection coefficient of the $l$-th target, including the propagation loss and the radar cross section (RCS), $\boldsymbol{\mathrm{a}}\left(\phi\right)\triangleq\left[1,e^{j\frac{2\pi}{\lambda}d\sin(\phi)},\dots,e^{j\frac{2\pi}{\lambda}(N-1)d\sin(\phi)}\right]^{T}\in\mathbb{C}^{N\times 1}$ is the steering vector of the DFRC antennas with $\lambda$ denoting the wavelength and $d$ denoting the transmit antenna spacing, and $\boldsymbol{W}$ is the AWGN, which is assumed to follow the CSCG distribution, i.e., $\boldsymbol{W}\sim\mathcal{CN}\left(0,\sigma_{\text{sen}}^{2}\textbf{I}\right)$. Moreover, we define $\boldsymbol{A}\triangleq\left[\boldsymbol{\mathrm{a}}\left(\phi_{1}\right),\cdots,\boldsymbol{\mathrm{a}}\left(\phi_{L}\right)\right]$ and $\boldsymbol{\Sigma}\triangleq\text{diag}\left(\alpha_{1},\cdots,\alpha_{L}\right)$.
@@
where $\dot{\boldsymbol{A}}_{l}\triangleq\dot{\boldsymbol{\mathrm{a}}}\left(\phi_{l}\right)\boldsymbol{\mathrm{a}}\left(\phi_{l}\right)^{T}+\boldsymbol{\mathrm{a}}\left(\phi_{l}\right)\dot{\boldsymbol{\mathrm{a}}}\left(\phi_{l}\right)^{T}$ with $\dot{\boldsymbol{\mathrm{a}}}\left(\phi_{l}\right)$ denoting the derivative of the steering vector, and $\dot{\boldsymbol{\mathrm{a}}}\left(\phi_{l}\right)$ can be expressed as
$\dot{\boldsymbol{\mathrm{a}}}\left(\phi\right)=$
$\left[0,j\frac{2\pi}{\lambda}d\text{cos}(\phi)\mathrm{a}(2),\cdots,j\frac{2\pi}{\lambda}(N-1)d\text{cos}(\phi)\mathrm{a}(N)\right]^{T},$ (11)
where $\mathrm{a}(n)$ denotes the $n$-th entry of $\dot{\boldsymbol{\mathrm{a}}}\left(\phi\right)$. The CRB are derived in Appendix A.
@@ III-B1 Approximation for Communication Rate
In order to approximate the tail probability (III-B1), we propose the following lemma (Lemma 1).
@@
Next, we use Lemma 1 to approximate (III-B1). First, we define $\boldsymbol{\Xi}_{\text{BU},k}\triangleq\gamma_{\text{BU},k}^{2}\textbf{I}$ and $\boldsymbol{\Xi}_{\text{BRU},k}\triangleq\gamma_{\text{BRU},k}^{2}\textbf{I}$. We have $\Delta\boldsymbol{h}_{\text{BU},k}=\gamma_{\text{BU},k}\boldsymbol{e}_{\text{BU},k}$ and $\Delta\boldsymbol{H}_{\text{BRU},k}=\gamma_{\text{BRU},k}\boldsymbol{E}_{\text{BRU},k}$, where $\boldsymbol{e}_{\text{BU},k}\sim\mathcal{CN}(0,\textbf{I})$ and $\boldsymbol{E}_{\text{BRU},k}\sim\mathcal{CN}(0,\textbf{I})$. Then, (III-B1) is reformulated as the following equation by substituting (5) and (6) into (III-B1), which is written as
@@
Proposition 2. The outage probability constraint (III-B1) can be safely approximated by Lemma 1 into the following second order cone (SOC) constraint

### v2

@@ Abstract
As a critical technology for next-generation communication networks, integrated sensing and communication (ISAC) aims to achieve the harmonious coexistence of communication and sensing. The degrees-of-freedom (DoF) of ISAC is limited due to multiple performance metrics used for communication and sensing. Reconfigurable Intelligent Surfaces (RIS) composed of metamaterials can enhance the DoF in the spatial domain of ISAC systems. However, the availability of perfect Channel State Information (CSI) is a prerequisite for the gain brought by RIS, which is not realistic in practical environments. Therefore, under the imperfect CSI condition, we propose a decomposition-based large deviation inequality approach to eliminate the impact of CSI error on communication rate and sensing Cramér-Rao bound (CRB). Then, an alternating optimization (AO) algorithm based on semi-definite relaxation (SDR) and gradient extrapolated majorization-maximization (GEMM) is proposed to solve the transmit beamforming and discrete RIS beamforming problems. We also analyze the complexity and convergence of the proposed algorithm. Simulation results show that the proposed algorithms can effectively eliminate the influence of CSI error and have good convergence performance. Notably, when CSI error exists, the gain brought by RIS will decrease with the increase of the number of RIS elements.
@@ I Introduction
Numerous emerging applications for future networks, including vehicle-to-everything (V2X), smart homes, and industrial Internet of thing (IoT), require high-quality communication performance and demand accurate sensing capabilities [1, 2]. The communication and sensing functionality can be simultaneously realized in a unified platform due to their similar hardware architecture and signal processing algorithms. Motivated by these considerations, research on integrated sensing and communication (ISAC) is currently underway to empower next-generation communication systems with sensing capabilities, such as mobile edge computing (MEC) [3], federated learning [4], and beamforming [5]. ISAC can realize dynamic spectrum sharing between sensing and communication, achieving harmonious co-existence between radar and communication systems [6].
Reconfigurable intelligent surface (RIS) is a planar surface consisting of a large number of low-cost and nearly passive reconfigurable elements that can be dynamically controlled to induce appropriate amplitude and phase shifts to incident signals [7, 8, 9, 10]. RIS has been identified as a promising technology to enhance the energy and spectral efficiency of communication systems and radar systems. Research on deploying RISs into communication systems mainly focuses on joint beamforming design [11, 12, 13], robust transmission design [14, 15], and energy-efficient design [16]. On the other hand, research on deploying RISs into radar systems focuses on enhancing sensing coverage in environments characterized by rich scattering, such as joint beamforming works [17, 18, 19], sensing architecture [20], and cellular sensing scheme via multiple RISs [21]. Additionally, RISs can be deployed in scenarios where infrastructures, such as the Global Positioning System (GPS), are insufficient to estimate the locations of UEs and scatterers. To be specific, theoretical bounds for localization have been investigated in [22], signal strength-based localization schemes have been proposed in [23], and time-of-arrival (TOA) and angle of departure (AOD) based methods have been investigated in [24, 25].
@@ I-A The Contribution of This Work
Our study focuses on robust beamforming for RIS-assisted ISAC systems. We propose a novel approximation method to approximate the probability constraints for the communication rate and the lower bound of the estimation accuracy, i.e., the Cramér-Rao bound (CRB). We then propose an alternating optimization (AO) algorithm to solve the robust beamforming problem, with two schemes for handling the non-convex discrete RIS phase shifts constraints. Our proposed robust beamforming method is demonstrated to approach the probability constraints tightly and handle the discrete RIS phase shifts effectively.
@@ II System Model
Figure 1: A RIS-assisted ISAC system.
The considered system is illustrated in Fig. 1, where an integrated sensing and communication (ISAC) BS with $N$ transmit antennas and $N_{\text{R}}$ receive antennas communicates with $K$ single-antenna user while sensing $L$ targets [^1] As illustrated by [20], the performance gain of the CRB of angle estimation from the reflected signal path of the passive RIS is minimal when the line-of-sight (LoS) signal paths between the BS and the targets exist. Therefore, we neglect the role of RIS in sensing. The robust transmission in non-line-of-sight (NLoS) scenarios will be discussed in our future work., with the aid of an $M$-element RIS. The transmit antennas and receive antennas are both uniform linear arrays (ULAs). The transmit and receive antennas at the BS are widely separated to suppress the leakage signals between them. We assume that the inter-distance of the transmit antennas, the receive antennas, and RIS elements are half-wavelength. The set of RIS elements, communication users, and targets are denoted by $\mathcal{M}$, $\mathcal{K}$, and $\mathcal{L}$. The phase shifts of the RIS elements are controlled by a RIS controller, which communicates wirelessly with the ISAC BS for cooperative transmission. The phase matrix of the RIS is represented by $\boldsymbol{\Theta}\triangleq\text{diag}\left(\theta_{1},\cdots,\theta_{M}\right)$. We assume an ideal RIS reflection magnitude model where each element is fixed at $|\theta_{m}|=1,\forall m\in\mathcal{M}$, as well as a practical RIS reflection phase shifts model where each $\theta_{m}$ can only take $d$ finite values, equally spaced in $[0,2\pi)$. The discrete set of phase shifts for each RIS element is given by
@@ II-A Communication Signal Model
where $\boldsymbol{h}_{\text{BU},k}\in\mathbb{C}^{N\times 1}$ is the channel vector between the BS and the $k$-th user, $\boldsymbol{h}_{\text{RU},k}\in\mathbb{C}^{M\times 1}$ is the channel vector between the RIS and the $k$-th user, and $\boldsymbol{H}_{\text{BR}}\in\mathbb{C}^{N\times M}$ is the channel matrix between the BS and the RIS. The transmit beamforming matrix of the ISAC BS is $\boldsymbol{S}=\left[\boldsymbol{s}_{1},\boldsymbol{s}_{2},\cdots,\boldsymbol{s}_{K}\right]\in\mathbb{C}^{N\times K}$. The transmit signal matrix is $\boldsymbol{X}=\left[\boldsymbol{x}_{1},\boldsymbol{x}_{2},\cdots,\boldsymbol{x}_{K}\right]^{T}\in\mathbb{C}^{K\times T}$, where $T$ is the number of signal samples. The signal vectors in $\boldsymbol{X}$ are assumed to be statistically independent, i.e., $\mathbb{E}\left[\boldsymbol{X}\boldsymbol{X}^{H}\right]=\textbf{I}_{K}$. Moreover, $\boldsymbol{z}_{k}$ is the additive white Gaussian noise (AWGN) at the $k$-th user, which is assumed to follow the circularly symmetric complex Gaussian (CSCG) distribution, i.e., $\boldsymbol{z}_{k}\sim\mathcal{CN}\left(0,\sigma_{\text{com},k}^{2}\textbf{I}\right)$ with $\sigma_{\text{com},k}^{2}$ denoting the variance of the AWGN.
Due to the passive characteristic of the RIS, the cascaded BS-RIS-user channel at the transmitter (CBRUT) is difficult to obtain. The direct channel state information at the transmitter (DCSIT) is also difficult to obtain in the complex wireless propagation environments. Hence, we assume the imperfect DCSIT and the imperfect CBRUT [^2] The channel estimation method in [36] can be used to estimated the imperfect DCSIT and the imperfect CBRUT.. The imperfect DBIUT can be expressed as
@@
where $\boldsymbol{H}_{\text{BRU},k}\triangleq\text{diag}\left(\boldsymbol{h}_{\text{RU},k}\right)^{H}\boldsymbol{H}_{\text{BR}}^{H}\in\mathbb{C}^{M\times N}$ is the cascaded BS-RIS-user channel. The CSI error, i.e., $\Delta\boldsymbol{h}_{\text{BU},k}$ and $\Delta\boldsymbol{H}_{\text{BRU},k}$ are all assumed to follow the CSCG distribution [^3] According to [36], the variances of the CSI error are only related to the transmit power of the signal during channel estimation, the received combining vectors, the covariance matrices of the received signals, and the noise variance. All of these parameters are known, so we assume that the variances of the CSI error are also known.. The CSI error can be expressed as
@@ II-B Sensing Signal Model
The received sensing signal at the ISAC BS is $\boldsymbol{Y}_{\text{sen}}$, which can be expressed as
@@
where $\boldsymbol{G}_{\text{sen},l}\triangleq\alpha_{l}\boldsymbol{\mathrm{b}}\left(\varphi_{l}\right)\boldsymbol{\mathrm{a}}\left(\phi_{l}\right)^{T}\in\mathbb{C}^{N\times N}$ denotes the sensing response matrix, $\alpha_{l}$ is the reflection coefficient of the $l$-th target, including the propagation loss and the radar cross section (RCS), $\phi_{l}$ and $\varphi_{l}$ are respectively the angle of departure (AoD) and the angle of arrival (AoA) of the $l$-th target, $\boldsymbol{\mathrm{a}}\left(\phi\right)\triangleq\left[1,e^{j\pi\sin(\phi)},\dots,e^{j\pi(N-1)\sin(\phi)}\right]^{T}\in\mathbb{C}^{N\times 1}$ is the steering vector of the ISAC transmit antennas, $\boldsymbol{\mathrm{b}}\left(\varphi\right)\triangleq\left[1,e^{j\pi\sin(\varphi)},\dots,e^{j\pi(N_{\text{R}}-1)\sin(\varphi)}\right]^{T}\in\mathbb{C}^{N_{\text{R}}\times 1}$ is the steering vector of the ISAC receive antennas, and $\boldsymbol{W}$ is the AWGN, which is assumed to follow the CSCG distribution, i.e., $\boldsymbol{W}\sim\mathcal{CN}\left(0,\sigma_{\text{sen}}^{2}\textbf{I}\right)$. Moreover, we define $\boldsymbol{A}\triangleq\left[\boldsymbol{\mathrm{a}}\left(\phi_{1}\right),\cdots,\boldsymbol{\mathrm{a}}\left(\phi_{L}\right)\right]$, $\boldsymbol{B}\triangleq\left[\boldsymbol{\mathrm{b}}\left(\varphi_{1}\right),\cdots,\boldsymbol{\mathrm{b}}\left(\varphi_{L}\right)\right]$, and $\boldsymbol{\Sigma}\triangleq\text{diag}\left(\alpha_{1},\cdots,\alpha_{L}\right)$.
It is noted that in Eq. (9), the echoes from $L$ targets are mixed together. We can use the MUSIC algorithm proposed in [46] to estimate the angles of multiple targets. This enables the feasibility of separately optimizing the CRB of angle estimation for each target. Additionally, we focus on the optimal achievable performance of angle estimation without concerning specific angle estimation algorithms.
@@
where $\dot{\boldsymbol{A}}_{l}\triangleq\dot{\boldsymbol{\mathrm{b}}}\left(\varphi_{l}\right)\boldsymbol{\mathrm{a}}\left(\phi_{l}\right)^{T}+\boldsymbol{\mathrm{b}}\left(\varphi_{l}\right)\dot{\boldsymbol{\mathrm{a}}}\left(\phi_{l}\right)^{T}$ with $\dot{\boldsymbol{\mathrm{a}}}\left(\phi_{l}\right)$ and $\dot{\boldsymbol{\mathrm{b}}}$ denoting the derivatives of the transmit steering vector and the receive steering vector, and they can be expressed as
$\dot{\boldsymbol{\mathrm{a}}}\left(\phi\right)=\left[0,j\pi\text{cos}(\phi)\mathrm{a}(2),\cdots,j\pi(N-1)\text{cos}(\phi)\mathrm{a}(N)\right]^{T},$ (11)
$\dot{\boldsymbol{\mathrm{b}}}\left(\varphi\right)=\left[0,j\pi\text{cos}(\varphi)\mathrm{b}(2),\cdots,j\pi(N_{\text{R}}-1)\text{cos}(\varphi)\mathrm{b}(N_{\text{R}})\right]^{T},$
where $\mathrm{a}(n)$ and $\mathrm{b}(n)$ denote the $n$-th entry of $\boldsymbol{\mathrm{a}}\left(\phi\right)$ and $\boldsymbol{\mathrm{b}}\left(\phi\right)$, respectively. The CRB are derived in Appendix A.
@@
Remark 2. The value of the CRB in (10) depends on the angle values. On one hand, we can minimize the CRB in the direction of interest, where the potential targets may exist. On the other hand, when the target is static or moves slowly, it is not necessary to update the beamforming matrices frequently, and the estimated angles previously are sufficient to characterize the CRB of angle estimation. Therefore, we assume that the angles of the targets are known.
@@ III-B1 Approximation for Communication Rate
$=\text{Pr}\left\{\boldsymbol{c}_{k}\left(\boldsymbol{\theta}\right)\left[\boldsymbol{s}_{k}\boldsymbol{s}_{k}^{H}-\left(2^{r_{k}}-1\right)\boldsymbol{S}_{-k}\boldsymbol{S}_{-k}^{H}\right]\boldsymbol{c}_{k}^{H}\left(\boldsymbol{\theta}\right)\geq 0\right\}$
@@
In order to approximate the tail probability (15), we propose the following lemma (Lemma 1).
@@
Next, we use Lemma 1 to approximate (15). First, we define $\boldsymbol{\Xi}_{\text{BU},k}\triangleq\gamma_{\text{BU},k}^{2}\textbf{I}$ and $\boldsymbol{\Xi}_{\text{BRU},k}\triangleq\gamma_{\text{BRU},k}^{2}\textbf{I}$. We have $\Delta\boldsymbol{h}_{\text{BU},k}=\gamma_{\text{BU},k}\boldsymbol{e}_{\text{BU},k}$ and $\Delta\boldsymbol{H}_{\text{BRU},k}=\gamma_{\text{BRU},k}\boldsymbol{E}_{\text{BRU},k}$, where $\boldsymbol{e}_{\text{BU},k}\sim\mathcal{CN}(0,\textbf{I})$ and $\boldsymbol{E}_{\text{BRU},k}\sim\mathcal{CN}(0,\textbf{I})$. Then, (15) is reformulated as the following equation by substituting (5) and (6) into (15), which is written as
@@
Proposition 2. The outage probability constraint (18) can be safely approximated by Lemma 1 into the following SOC constraint